If a basketball player has a vertical leap of about 35 inches, what is his hang tome?Hang time function is v=48T^2
Answer by checkley71(8403)

v=48t^2
35=48t^2
t^2=35/48
t^2=.7292
t=.854 (seconds) is the hang time.